# Can Helicity Modulus Be Defined For Boundary Conditions With Finite Twist?

### v3

@@ Abstract
We study the response of a two-dimensional classical XY model to a finite (non-infinitesimal) twist of the boundary conditions. We use Monte Carlo simulations to evaluate the free energy difference between periodic and twisted-periodic boundary conditions and find deviations from the expected quadratic dependence on the twist angle. Consequently, the helicity modulus (spin-stiffness) shows a non-trivial dependence on the twist angle. We show that the deviation from the expected behavior arises because of the mixing of states with opposite chirality which leads to an additional entropy contribution in the quasi-long-range ordered phase. We give an improved prescription for the numerical evaluation of the helicity modulus, and we discuss the spontaneous breaking of the chiral symmetry for anti-periodic boundary conditions. We also discuss applications to discrete spin systems and some experimental scenarios where boundary conditions with finite twist are necessary.
@@ I Introduction
Even in system in which the boundary conditions do not affect the bulk behavior, they can have interesting and observable consequences. For example, thermodynamic Casimir forces [9], which arise due to the confinement of critical order parameter fluctuations, were found to be affected by the boundary conditions [10, 11, 12, 13, 14]. Moreover, the response of a system to a change in boundary conditions can be employed as a tool to measure (bulk) equilibrium properties.
In this paper, we reconsider the seemingly simple but surprisingly complex effects of a finite (non-infinitesimal) twist of the boundary conditions on a ferromagnetic XY system. Arguments put forward by Brown and Ciftan [15] in the context of the three-dimensional Heisenberg model suggest that the mixing of states with different chirality plays an important role and affects the observed helicity modulus. However, a quantitative understanding of this effect and its origins has not been achieved yet. We therefore study this question in detail for the two-dimensional classical XY model by means of large-scale Monte Carlo simulations. We find that the free energy cost of a non-infinitesimal twist in the boundary condition in the quasi long-range ordered (QLRO) low-temperature phase deviates from the expected quadratic dependence on the twist angle. Mixing of states with opposite chirality provides an extra entropy contribution which takes the value $\ln(2)$ (in units in which $k_{B}=1$) for a $\pi$-twist. Beyond their intrinsic interest, our results potentially apply to experiments aimed at detecting the Berezinskii–Kosterlitz–Thouless (BKT) transition [16, 17]. They are also important for systems with discrete $Z(N)$ (clock) order parameter symmetry where any twist in the boundary conditions is necessarily non-infinitesimal. In addition, our findings enable novel numerical algorithms for computing the helicity modulus in simulations.
@@ II The Model
The definition (4) has the advantage that the second derivative of the free energy can be expressed in terms of appropriate correlation functions of the untwisted system. This leads to the formula [22]
$\left(\frac{\partial^{2}F}{\partial\Theta^{2}}\right)_{\Theta=0}=\frac{1}{L^{2-d}}\sum_{<ij>}J\left\langle\cos(\phi_{i}-\phi_{j})\right\rangle(x_{i}-x_{j})^{2}$
$-\frac{\beta}{L^{2-d}}\left\langle\left\{\sum_{<ij>}J\sin(\phi_{i}-\phi_{j})(x_{i}-x_{j})\right\}^{2}\right\rangle$ (5)
where $x_{i}$ is the coordinate of site $i$ in the twisted direction, and delimited-⟨$\langle$..$\rangle$ … ⟩ denotes the thermodynamic average evaluated at $\Theta=0$. This formula allows the evaluation of the helicity modulus without actually having to apply twisted boundary conditions.
As the paramagnetic phase is insensitive to the boundary conditions, the free energy difference decays exponentially with system size, $\Delta F\sim e^{-L/\xi}$ , where $\xi$ is the correlation length. This implies $\rho_{s}=0$ in the paramagnetic phase in the thermodynamic limit. In contrast, $\Delta F$ is expected to scale as $L^{d-2}$ in an ordered or quasi long-range ordered phase, and the helicity modulus is finite. In two dimensions, $\rho_{s}$ is known to have an universal jump at the BKT phase transition. For $d>2$, $\rho_{s}$ vanishes continuously at $T_{c}$, governed by critical behavior of the phase transition. In the rest of the paper, we focus on two dimensions, but we will comment on higher dimensions in the concluding section.
@@ IV.1 Free energy response to finite twist
We now turn to our results for the response of the two-dimensional XY model to various twists in the boundary conditions. Figure 1 shows the helicity modulus $\rho_{s}(\Theta)$ as a function of temperature for three different cases, viz., an infinitesimal twist, a twist of $\Theta=\pi/6$, and a twist of $\Theta=\pi$.
Figure 1: Helicity modulus $\rho_{s}(\Theta)$ as a function of temperature $T$ for different system sizes $L$. The filled symbols show $\rho_{s}$ for a twist of $\Theta=\pi/6$, obtained from the definition eq. (3) using the free energies eq. (6). The dotted lines show $\rho_{s0}$ for an infinitesimal twist measured in the untwisted system via eq. (5). The helicity modulus values for these two cases agree within their statistical errors (which are comparable to or smaller than the symbol sizes). The straight solid line corresponds to $\rho_{s}=2T/\pi$, it intersects the stiffness curves at temperatures $T^{*}(L)$. The critical temperature $T_{c}$ is obtained by extrapolating $T^{*}(L)$ to the thermodynamic limit according to eq. (7). This extrapolation gives $T_{c}=0.891(7)$. In contrast, the helicity modulus values for a twist of $\Theta=\pi$, shown as dash-dotted line for $L=40$, are significantly lower than the other data.
For the infinitesimal twist, $\rho_{s}$ is measured in the untwisted system via eq. (5). For the finite twists, $\rho_{s}$ is obtained from the free energy difference between simulations with twisted and untwisted boundary conditions, as explained in Sec. III. The resulting $\rho_{s}$ values for the infinitesimal twist and $\Theta=\pi/6$ agree within their statistical errors, giving us additional confidence in our numerical approach.
@@
We also studied a twist of $\Theta=\pi$. Unexpectedly, the helicity modulus values resulting from eq. (3) in this case are significantly below the values for smaller $\Theta$, as shown in Fig. 1. It is worth emphasizing that this happens even though the layer-to-layer twist $\Theta/L$ remains small compared to unity, justifying the expansion that leads to eq. (2). What is the reason for this surprising discrepancy? Arguments put forward in Ref. [15] suggest that the chirality of the twist plays an important role.
So far we have considered thermodynamic states that fulfill the global twist $\Theta$ in, say, clockwise direction by introducing an average clockwise twist of $\Theta/L$ between neighboring layers. However, the same global twist can be achieved by a counter-clockwise twist with local angle of $(2\pi-\Theta)/L$ between consecutive layers. At low temperatures, the additional free energy for a twist in the “wrong” direction is much larger than $k_{B}T$. Therefore, this state is exponentially suppressed. Local twist angles corresponding to higher winding numbers can also be ruled out using the same argument. However, with increasing temperature, states of both chiralities (and higher winding numbers) will be mixed, leading to an extra entropic contribution to the free energy. Importantly, $\Theta=\pi$ is a special case that allows the mixing of opposite chiralities even at $T=0$. Let us now quantitatively study the mixing of the states as twist, temperature and system size is varied.
@@
Figure 2: Free energy difference $\Delta F$ vs. squared twist angle $\Theta^{2}$ for size $L=40$ and several temperatures close to $T_{c}$. The dotted line shows a fit of $\Delta F$ for $T=0.82$ to a quadratic $\Theta$ dependence in the range $0\leq\Theta^{2}\leq(2\pi/3)^{2}$. $\Delta F$ at $T_{c}\approx 0.893$, obtained using eq. (8) is marked by the star symbol.
The figure shows that $\Delta F$ follows the quadratic $\Theta$ dependence up to about $\Theta=2\pi/3$ but significant deviations are observed for larger twist angles. The special case of anti-periodic boundary conditions ($\Theta=\pi$) right at the BKT transition temperature was already studied by Hasenbusch [34, 35]. He derived an expression for the ratio of the partition functions with periodic and antiperiodic boundary conditions in the two-dimensional XY model. It includes the leading finite-size corrections and reads
@@
We attribute the deviation of $\Delta F$ from the quadratic $\Theta$ dependence to the mixing of states of opposite chirality which becomes more pronounced as $\Theta$ approaches $\pi$. To quantify this mechanism, we determine the excess free energy due to the mixing. To this end, we first fit a quadratic function to $\Delta F(\Theta)$ in the range $30\leq\Theta\leq 2\pi/$ 3 (separately for each system size). Denoting the fit function by $\Delta F_{\mathrm{fit}}(L,T,\Theta)$, we define the excess free energy $\delta\Delta F$ at twist angle $\pi$ as the difference between $\Delta F_{\mathrm{fit}}(L,T,\pi)$ and the $\Delta F_{\mathrm{MC}}(L,T,\pi)$ obtained from MC,
@@
Figure 3: Excess free energy $\delta\Delta F(L,T,\pi)$ defined in eq. (9) for a twist angle of $\pi$ as a function of temperature $T$ for different system sizes. The solid line represents the function $T\ln(2)$.
Figure 3 shows the excess free energy $\delta\Delta F$ for twist angle $\pi$ as a function of temperature $T$ for various system sizes. For temperatures $T>T_{c}$, in the disordered phase, $\delta\Delta F$ decreases with increasing system size and appears to vanish in the thermodynamic limit. For temperatures $T<T_{c}$, in contrast, $\delta\Delta F$ approaches a nonzero limit for $L\to\infty$. This infinite-system limit of $\delta\Delta F$ follows the function $T\ln(2)$ as temperature goes to zero  [^1] The small deviations of $\delta\Delta F$ from $T\ln(2)$ for smaller system sizes persist to the lowest temperatures. They can be explained quantitatively by contributions from the quartic terms in $(\Theta/L)$ that arise in the expansion of the zero-temperature energy difference $\Delta U=U(\Theta)-U(0)=-JL^{2}(1-\cos(\Theta/L))$..
@@ IV.2 Breaking of the chiral symmetry
In Sec. IV.1, we have established that, for an imposed twist $\Theta=\pi$, the thermodynamic ensemble contains an equal mixture of states with clockwise and counter-clockwise local twists. Now we explore the question whether a single given macroscopic system fluctuates between clockwise and counter-clockwise orientations or whether it spontaneously breaks the chiral symmetry below $T_{c}$ by freezing into one orientation. As the classical XY Hamiltonian does not contain any dynamic terms, the answer to this question will depend on the assumed dynamics of the system. Here, we focus on dissipative local model-A dynamics according to the Hohenberg-Halperin classification [38], but the results should hold qualitatively for other local dynamics as well.
To analyze chiral fluctuations, we define the bulk chirality $h$ via the vector product between nearest neighbor spins along the direction of the twist. Assuming a twist in the $x$-direction, $h$ is given by
@@
The results of this section demonstrate that the energy required to flip the bulk chirality diverges with system size in the thermodynamic limit. The divergence is only logarithmic in $L$ in contrast to the case of a domain wall whose energy would diverge linearly in $L$ in two dimensions. Nonetheless, it implies that a macroscopic system will not fluctuate between states of opposite bulk chirality during its time evolution (for local dynamics). In other words, the chiral symmetry is the spontaneously broken below the BKT transition.

### v4

# Helicity modulus and chiral symmetry breaking for boundary conditions with finite twist

@@ Abstract
We study the response of a two-dimensional classical XY model to a finite (non-infinitesimal) twist of the boundary conditions. We use Monte Carlo simulations to evaluate the free energy difference between periodic and twisted-periodic boundary conditions and find deviations from the expected quadratic dependence on the twist angle. Consequently, the helicity modulus (spin-stiffness) shows a non-trivial dependence on the twist angle. We show that the deviation from the expected behavior arises because of the mixing of states with opposite chirality which leads to an additional entropy contribution in the quasi-long-range ordered phase. We give an improved prescription for the numerical evaluation of the helicity modulus for a finite twist, and we discuss the spontaneous breaking of the chiral symmetry for the anti-periodic boundary conditions. We also discuss applications to discrete spin systems and some experimental scenarios where boundary conditions with finite twist are necessary.
@@ I Introduction
Even in system in which the boundary conditions do not affect the bulk behavior, they can have interesting and observable consequences. For example, thermodynamic Casimir forces [9], which arise due to the confinement of critical order parameter fluctuations, were found to be affected by the boundary conditions [10, 11, 12, 13, 14] (for a recent review of exact results see Ref. [15]). An analogous force, the Helmholtz force, arises in the canonical ensemble and also shows a strong dependence on the boundary conditions [16]. Moreover, the response of a system to a change in boundary conditions can be employed as a tool to measure (bulk) equilibrium properties.
In this paper, we reconsider the seemingly simple but surprisingly complex effects of a finite (non-infinitesimal) twist of the boundary conditions on a ferromagnetic XY system. Brown and Ciftan [17] suggested, in the context of the three-dimensional Heisenberg model, that the mixing of states with differing chirality plays an important role and affects the observed helicity modulus. However, a quantitative understanding of this effect and its origins has not been achieved yet. We therefore study this question in detail for the two-dimensional classical XY model by means of large-scale Monte Carlo simulations.
We find that the free energy cost of a non-infinitesimal twist in the boundary condition in the quasi long-range ordered (QLRO) low-temperature phase deviates from the expected quadratic dependence on the twist angle. Mixing of states with opposite chirality provides an extra entropy contribution which takes the value $\ln(2)$ (in units in which $k_{B}=1$) for a $\pi$-twist. Beyond their intrinsic interest, our results potentially apply to experiments aimed at detecting the Berezinskii–Kosterlitz–Thouless (BKT) transition [18, 19]. They are also important for systems with discrete $Z(N)$ (clock) order parameter symmetry where any twist in the boundary conditions is necessarily non-infinitesimal. In addition, our findings enable novel numerical algorithms for computing the helicity modulus in simulations with a finite twist.
@@ II The Model
The definition (4) has the advantage that the second derivative of the free energy can be expressed in terms of appropriate correlation functions of the untwisted system. This leads to the formula [24, 25]
$\left(\frac{\partial^{2}F}{\partial\Theta^{2}}\right)_{\Theta=0}=\frac{1}{L^{2}}\sum_{<ij>}J\left\langle\cos(\phi_{i}-\phi_{j})\right\rangle(x_{i}-x_{j})^{2}$
$-\frac{\beta}{L^{2}}\left\langle\left\{\sum_{<ij>}J\sin(\phi_{i}-\phi_{j})(x_{i}-x_{j})\right\}^{2}\right\rangle$ (5)
where $x_{i}$ is the coordinate of site $i$ in the twisted direction, and delimited-⟨$\langle$..$\rangle$ … ⟩ denotes the thermodynamic average evaluated at $\Theta=0$. For a derivation of Eq. (5), see Appendix A. This formula allows the evaluation of the helicity modulus without actually having to apply twisted boundary conditions.
As the paramagnetic phase is insensitive to the boundary conditions, the free energy difference decays exponentially with system size, $\Delta F\sim e^{-L/\xi}$ , where $\xi$ is the correlation length. This implies $\rho_{s}=0$ in the paramagnetic phase in the thermodynamic limit. In contrast, $\Delta F$ is expected to scale as $L^{d-2}$ in an ordered or quasi long-range ordered phase, and the helicity modulus is finite. In two dimensions, $\rho_{s}$ is known to have an universal jump at the BKT phase transition. For $d>2$, $\rho_{s}$ vanishes continuously at $T_{c}$, governed by the critical behavior of the phase transition. In the rest of the paper, we focus on two dimensions, but we will comment on higher dimensions in the concluding section.
@@ IV.1 Free energy response to finite twist
We now turn to our results for the response of the two-dimensional XY model to various twists in the boundary conditions. Figure 1(a) shows the helicity modulus $\rho_{s}(\Theta)$ as a function of temperature for an infinitesimal twist as well as a twist of $\Theta=\pi/6$.
Figure 1: (a) Helicity modulus $\rho_{s}(\Theta)$ as a function of temperature $T$ for different system sizes $L$. The filled symbols show $\rho_{s}$ for a twist of $\Theta=\pi/6$, obtained via Eqs. (3) and (6). The dotted lines show $\rho_{s0}$ for an infinitesimal twist measured in the untwisted system via Eq. (5). The straight solid line corresponds to $\rho_{s}=2T/\pi$, it intersects the stiffness curves at temperatures $T^{*}(L)$. The critical temperature $T_{c}$ is obtained by extrapolating $T^{*}(L)$ to the thermodynamic limit according to Eq. (7). This extrapolation gives $T_{c}=0.891(7)$. In contrast, the helicity modulus values for a twist of $\Theta=\pi$, shown as dash-dotted line for $L=56$, are significantly lower than the other data. (b) Helicity modulus $\rho_{s}(\Theta)$ as a function of temperature $T$ at fixed $L=40$ for different twist angles $\Theta$. The statistical errors in both panels are comparable to or smaller than the symbol sizes.
For the infinitesimal twist, $\rho_{s}$ is measured in the untwisted system via eq. (5). For a finite twist, $\rho_{s}$ is obtained from the free energy difference between simulations with twisted and untwisted boundary conditions, as explained in Sec. III. The resulting $\rho_{s}$ values for the infinitesimal twist and $\Theta=\pi/6$ agree within their statistical errors, giving us additional confidence in our numerical approach.
@@
Figure 1(a) also shows data for $\Theta=\pi$. Unexpectedly, the helicity modulus values resulting from eq. (3) in this case are significantly below the values for smaller $\Theta$. This is further illustrated in Fig. 1(b) which compares the helicity modulus for different twist angles $\Theta$. Whereas the data for $\Theta=0,\pi/6$, and $\pi/2$ all agree within their error bars, some deviations appear for $\Theta=5\pi/6$ close to the BKT transition. They become more pronounced for $\Theta=\pi$ and persist in the entire quasi-long-range ordered phase. It is worth emphasizing that this happens even though the layer-to-layer twist $\Theta/L$ remains small compared to unity, justifying the expansion that leads to eq. (2). What is the reason for this surprising discrepancy? Arguments put forward in Ref. [17] suggest that the chirality of the twist plays an important role.
So far we have considered thermodynamic states that fulfill the global twist $\Theta$ in, say, the clockwise direction by introducing an average clockwise twist of $\Theta/L$ between neighboring layers. However, the same global twist can be achieved by a counter-clockwise twist with local angle of $(2\pi-\Theta)/L$ between consecutive layers. At low temperatures, the additional free energy for a twist in the “wrong” direction is much larger than $k_{B}T$. Therefore, this state is exponentially suppressed. Local twist angles corresponding to higher winding numbers can also be ruled out using the same argument. However, with increasing temperature, states of both chiralities (and higher winding numbers) will be mixed, leading to an extra entropic contribution to the free energy. Importantly, $\Theta=\pi$ is a special case that allows the mixing of opposite chiralities even at $T=0$. Let us now quantitatively study the mixing of the states as twist, temperature and system size is varied.
@@
Figure 2: Free energy difference $\Delta F$ between systems with periodic and twisted-periodic boundary conditions vs. squared twist angle $\Theta^{2}$ for size $L=40$ and several temperatures close to $T_{c}$. The dashed line shows a fit of $\Delta F$ for $T=0.82$ to a quadratic $\Theta$ dependence in the range $0\leq\Theta^{2}\leq(2\pi/3)^{2}$. $\Delta F$ at $T_{c}\approx 0.893$, obtained using eq. (8) is marked by the star symbol. Statistical errors are comparable to the symbol size or smaller.
The figure shows that $\Delta F$ follows the quadratic $\Theta$ dependence up to about $\Theta=2\pi/3$ within the statistical errors of the data, but significant deviations are observed for larger twist angles.
The special case of anti-periodic boundary conditions ($\Theta=\pi$) right at the BKT transition temperature was already studied by Hasenbusch [37, 38]. He derived an expression for the ratio of the partition functions with periodic and antiperiodic boundary conditions in the two-dimensional XY model. It includes the leading finite-size corrections and reads
@@
We attribute the deviation of $\Delta F$ from the quadratic $\Theta$ dependence to the mixing of states of opposite chirality which becomes more pronounced as $\Theta$ approaches $\pi$. The extent of the mixing can be estimated by studying a simple two-state model consisting of a state with clockwise (CW) chirality of the twist and a state with counter-clockwise (CCW) chirality. At low temperatures and large $L$, the energy associated with the CW state reads $U_{\mathrm{CW}}=-JL^{2}[1+\cos(\Theta/L)]\approx-2JL^{2}+J\Theta^{2}/2$, whereas the energy of the CCW state is given by $U_{\mathrm{CCW}}=-JL^{2}[1+\cos((2\pi-\Theta)/L)]\approx-2JL^{2}+J(2\pi-\Theta)^{2}/2$. At higher temperatures, fluctuations about the perfect spin-wave states become important. Their effect can be approximately captured by replacing $J$ in the above formulas for $U_{\mathrm{CW}}$ and $U_{\mathrm{CCW}}$ by the (renormalized) helicity modulus $\rho_{s}(T)$. The corresponding canonical probabilities of CW and CCW states read
$p_{\mathrm{CW}}=\frac{1}{Z}e^{-\frac{\rho_{s}\Theta^{2}}{2T}}~{},\quad p_{\mathrm{CCW}}=\frac{1}{Z}e^{-\frac{\rho_{s}(2\pi-\Theta)^{2}}{2T}}$ (9)
with
$Z=e^{-\frac{\rho_{s}\Theta^{2}}{2T}}~{}+~{}e^{-\frac{\rho_{s}(2\pi-\Theta)^{2}}{2T}}~{}.$ (10)
As discussed earlier, this implies that the CCW state is exponentially suppressed as $T\to 0$ for any $\Theta<\pi$ whereas both states contribute equally at all temperatures below $T_{c}$ for $\Theta=\pi$.
The helicity modulus right at the BKT transition temperature $T_{c}$ is known from the Kosterlitz-Nelson relation, $\rho_{s}(T_{c})=\frac{2T_{c}}{\pi}$. Inserting this into eq. (9) gives the canonical probabilities shown in the inset of Fig. 3.
Figure 3: Free energy difference $\Delta F$ at the BKT transition temperature $T_{c}$ between systems with periodic and twisted-periodic boundary conditions, as evaluated using the two-state model (dashed line). The solid line corresponds to a quadratic fit in the range $0\leq\Theta^{2}\leq(2\pi/3)^{2}$. The value of $\Delta F$ from eq. (8) in the thermodynamic limit is marked by the star symbol. The canonical probabilities of the two states are shown in the inset as functions of $\Theta$.
The contribution of the CCW state to the mixture increases with $\Theta$ and becomes significant for $\Theta\gtrapprox 2\pi/3$. The free energy difference $\Delta F$ between systems with periodic and twisted-periodic boundary conditions can also be computed within the two-state model,
$\Delta F=\langle\Delta U\rangle-TS$ (11)
where $S$ is the Von-Neumann entropy. The main panel of Fig. 3 shows $\Delta F$ as a function of $\Theta$ at the BKT transition temperature. The deviation of $\Delta F$ from the quadratic dependence on $\Theta$ resembles the corresponding Monte Carlo result in Fig. 2. Moreover, $\Delta F$ from Eq. (8) in the thermodynamic limit $L=\infty$ (marked by the star symbol) agrees well with the model results. The simple two-state system thus captures the important features present in Fig. 2. This supports the notion that the mixing of states with opposite chiralities leads to the deviation of $\Delta F$ from the quadratic $\Theta$ dependence.
We now move beyond the two-state model and quantify the mixing of chiralities in the Monte Carlo data. To determine the excess free energy due to the mixing, we first fit a quadratic function to $\Delta F(\Theta)$ in the range $30\leq\Theta\leq 2\pi/$ 3 (separately for each system size). Denoting the fit function by $\Delta F_{\mathrm{fit}}(L,T,\Theta)$, we define the excess free energy $\delta\Delta F$ at twist angle $\pi$ as the difference between $\Delta F_{\mathrm{fit}}(L,T,\pi)$ and the $\Delta F_{\mathrm{MC}}(L,T,\pi)$ obtained from MC,
@@ IV.2 Breaking of the chiral symmetry
In Sec. IV.1, we have established that, for an imposed twist $\Theta=\pi$, the thermodynamic ensemble contains an equal mixture of states with clockwise and counter-clockwise local twists. Now we explore the question whether a single given macroscopic system fluctuates between clockwise and counter-clockwise orientations or whether it spontaneously breaks the chiral symmetry below $T_{c}$ by freezing into one orientation. As the classical XY Hamiltonian does not contain any dynamic terms, the answer to this question will depend on the assumed dynamics of the system. Here, we focus on dissipative local model-A dynamics according to the Hohenberg-Halperin classification [40], i.e., a purely relaxational dynamics without conservation laws. However, the results should hold qualitatively for other local dynamics as well.
To analyze chiral fluctuations, we define the bulk chirality $h$ via the vector product between nearest-neighbor spins along the direction of the twist. Assuming a twist in the $x$-direction, $h$ is given by
@@
The results of this section demonstrate that the energy required to flip the bulk chirality diverges with system size in the thermodynamic limit. The divergence is only logarithmic in $L$ in contrast to the case of a domain wall whose energy would diverge linearly in $L$ in two dimensions. Nonetheless, it implies that a macroscopic system will not fluctuate between states of opposite bulk chirality during its time evolution (for local dynamics). In other words, the chiral symmetry is spontaneously broken below the BKT transition.